Number of solutions of the equation tan x + sec x = 2 cos x, lying in the interval [0, 2p] is

Number of solutions of the equation  tan x + sec x = 2 cos x, lying in the interval [0, 2p] is

Grade:12

1 Answers

bharat bajaj IIT Delhi
9 years ago
tan x + sec x = 2cosx
sin x + 1 = 2cos2x
sin x + 1 = 2 (1 - sin2x)
2 sin2x + sin x - 1 = 0
Hence, we get
sin x = - 1, or sin x = 1/2
sin x = -1 orcosx = 0, which is not possible
? sin x = 1/2
x = p or x = p - (p / 6), in the interval [0, 2p]
x = p or x = 5p / 6
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty